Need the answers fast.

The joint probability mass function of (X,Y) is p(x, y) = k(2x + 3y) ; x = 0,1,2; y = 1,2,3. (i)
Find the marginal distributions. (ii) Find P(X = xi / Y = 2) (iii) Find P[X +Y > 3] .
1
Expert's answer
2020-08-09T18:31:33-0400
"p(x, y) = k(2x + 3y)"

for "X=0,1,2" and "Y = 1,2,3"


Solutions i)


Marginal distribution of X


"p(x) =\sum_{\ {\forall Y\\}} p(x,y)""= \sum_{\ {1<y<3\\}} k(2x + 3y)"

"= k(2x + 3) + k(2x + 6) + k(2x + 9)"

"=6kx + 18k"

Answer: p(x) = 6kx + 18k


Marginal distribution of Y


"p(y) =\sum_{\ {\forall X \\}} p(x,y)"

"= \sum_{\ {0<x<2\\}} k(2x + 3y)"

"= k(3y) + k(2 + 3y) + k(4 + 3y)"

"=9ky + 6k"

Answer: p(y) = 9ky + 6k



"\sum_{\ {\forall X \\}} p(x) = \sum_{\ {\forall Y \\}} p(Y) = 1"


"\sum_{\ {\forall X \\}} p(x) = \sum_{\ {0<x<2 \\}} 6kx + 18k"

"1 = (18k) + (6k + 18k) + (12k + 18k)"

"1 = 72k"

"=\frac{1}{72}"

"p(x,y) = \frac{2x+3y}{72}"

"p(x) = \frac{x}{12} + \frac{1}{4}"


"p(y) = \frac{y}{8} + \frac{1}{12}"

Solution ii) P(X = x / Y = 2)



"p(x,2) = \frac{2x+6}{72}"

Answer: ( 2 x + 6 ) / 72


Soultion iii) P[X +Y > 3]



"P[X +Y > 3] = p(X=2 , Y=2) or P(X=2 , Y=3) or P(X=1 , Y=3)"

"=\frac{4+6}{72} + \frac{4+9}{72} + \frac{2+9}{72} = \frac{17}{36}"

Answer: 17 / 36